OpenStudy (anonymous):

I click on Rbrass33's question about setting up an integral, and I get nothing but blank space. Something's terribly amiss.

OpenStudy (anonymous):

Is there supposed to be an integrand other than 1, or is the integral used to find the volume of the region? Is the integral supposed to be a single integral of is it supposed to be a multiple integral?

OpenStudy (anonymous):

hey sorry i don't know what wrong

OpenStudy (anonymous):

but i can post it here :) need help setting up integral. Let W be the region inside the cone z =sqrt(x^2 + y^2) and below the plane z = 1

OpenStudy (anonymous):

Looking for volume of region?

OpenStudy (anonymous):

yupp

OpenStudy (anonymous):

i was trying to use cylindrical coordinates and i had r is from 0 to 1, theta is from 0 to 2pi and z is from 0 to 1

OpenStudy (anonymous):

We can set up the integral using either the disk method or the cylindrical shell method.

OpenStudy (anonymous):

okay which one is easier

OpenStudy (anonymous):

Looks like disk method is easier. Dependent variable z goes from 0 1o 1, and we can partition it. In a particular subinterval of length dz is some particular z.

OpenStudy (anonymous):

The disk has radius z, so its volume dV is pr*z^2 dz. Since z goes from 0 to 1, \[V =\pi \int\limits_{0}^{1} z ^{2}dz.\]

OpenStudy (anonymous):

so u don't need to set up a triple integral to find the volume?

OpenStudy (anonymous):

What for? Keep it simple.

OpenStudy (anonymous):

lol i like that

OpenStudy (anonymous):

Even the cylindrical shell would be easier than the triple integral method. We could do it if we had to, though.

OpenStudy (anonymous):

thanks for the help :)

OpenStudy (anonymous):

i was just looking over the problem again and it says to evaluate the triple integral (1+sqrt(x^2+y^2) dV bound by W= the region inside the cone z =sqrt(x^2 + y^2) and below the plane z = 1

OpenStudy (anonymous):

would that still be equal to the volume= integral pi z^2 dz?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Yhe integral from 0 to 1, that is.